# Reasoning Limitations of Multimodal Large Language Models. A case study of Bongard Problems

### v1

@@ Appendix B Evaluation of MLLMs answers
Open-ended characteristics of BPs stemming from a textual form of an answer, and the number of considered models (8), generation strategies (7), datasets (4), and BP instances per dataset (60 in Bongard-RWR and $100$ in the remaining cases) require the use of an automated NLP-based evaluation of the model’s answers. For this task we employed MLLMs with a specially designed prompt. The initial version of the evaluation prompt (see Prompt B) was intentionally relatively simple – a model received an answer to be evaluated as well as the ground-truth labels, and was requested to output a binary yes/no answer. This prompt formulation turned out to be too simplistic. While the level of agreement between all models was relatively high ($87\%$ of responses were rated unanimously by all models), as illustrated in Fig. 4(a), manual inspection of the selected answers revealed that the assessment was generally too optimistic and relatively many evaluations wrongly pointed to correct answers.
@@
Prompt 14: Initial prompt used in MLLM answer evaluation. We focused on its clarity and simplicity. You are a logic module designed to provide accurate answers. In a Bongard Problem the objective is to spot the difference between the contents of images located on the two opposite sides of the problem. You are given correct labels of these sides and must decide whether the answer provided by the user is correct and matches with those labels. Answer with ’OK’ or ’WRONG’. LEFT SIDE LABEL: <left_label> RIGHT SIDE LABEL: <right_label> USER ANSWER: <model_answer>
B.1 Evaluation prompt optimization
Due to the above evaluation disagreement, we made an attempt to optimize the prompt based on the GPT-4o solutions for the additional $20$ BPs (#101 – #120) that were not used in the main experiments. First, following the few-shot prompting technique, we expanded the prompt to include two examples showing a possible logical difference between correct and incorrect answers. Furthermore, we added a sentence which requested a strict logical compliance with the provided labels. However, this refinement appeared to be too strong, as $2$ (out of $4$) models didn’t evaluate any of the solutions as correct.
To impose some flexibility, we changed the word strictly to logically, but this resulted in an increased rate of false-positive evaluations. Finally, we combined these two prompts, obtaining the outcome closest to the manual (our human) evaluation. The final version of our evaluation prompt is listed in Prompt B.1. Additionally, we attempted to attach the image of the evaluated BP instance to each version of our prompt, but this actually confused the models rather than improving their results, so we ultimately abandoned this option and stuck with the fully text-based prompt.
Although the consistency of results regarded as the number of unanonimous assessments stayed at the same level ($87\%$) (see Fig. 4(b)), the number of answers rated as correct significantly decreased, which was in accordance with our random manual verification.
@@ B.1 Evaluation prompt optimization
Prompt 15: The final version of the evaluation prompt. It is enriched with the few-shot prompting technique and imposes a logical compliance with provided labels. You are a logic module designed to provide accurate answers. In a Bongard Problem the objective is to spot the difference between the contents of images located on the two opposite sides of the problem. You are given correct labels of these sides and must decide whether the answer provided by the user is correct and matches with those labels. Answer with ’OK’ or ’WRONG’. The user’s answer has to strictly logically match the labels, as shown in examples. FIRST EXAMPLE: LEFT SIDE LABEL: All shapes are small. RIGHT SIDE LABEL: All shapes are big. USER ANSWER: On the left side, one of the shapes is small. On the right side, all of the shapes are big. EVALUATION: WRONG END OF FIRST EXAMPLE. SECOND EXAMPLE: LEFT SIDE LABEL: All shapes are small. RIGHT SIDE LABEL: All shapes are big. USER ANSWER: On the left side, all of the shapes are small. On the right side, all of the shapes are big. EVALUATION: OK END OF SECOND EXAMPLE. LEFT SIDE LABEL: <left_label> RIGHT SIDE LABEL: <right_label> USER ANSWER: <model_answer>
B.2 Manual verification of the evaluation performance of MLLMs
@@
In order to finally assess the efficacy of Prompt B.1 we manually checked the models’ evaluation performance on the $100$ BPs solved by GPT-4o using the Descriptive strategy. As shown in Table 1 in the main paper, all proprietary models achieved better scores on incorrect labels classification. For this reason, we decided to manually verify only the problems evaluated as correct by at least one MLLM, assuming that those incorrect are generally evaluated properly. The comparison between the evaluation performance of the initial and final prompts and our manual evaluation is presented in Table 2. All models denotes evaluations where a solution is marked as correct only if all models evaluate it as correct. Similarly, any model refers to the cases where a solution is marked as correct if at least one model evaluates it as correct. Voting refers to the hard-voting scheme described in section B.1. It is important to observe that the chosen voting evaluation method differed from the manual evaluation in only one specific problem, which is depicted in Fig. 5.
@@
In addition, we checked the performance of our enhanced evaluation prompt on $20$ new, not used in other experiments, manually evaluated Bongard-OpenWorld problems solved by GPT-4o using the Descriptive strategy. Again, the use of Prompt B.1 visibly increased the consensus with manual evaluation (see Table 3). The difference between our manual evaluation and the voting scheme occurred only in 2 problem solutions whose correctness is disputable (see Fig. 6).
@@
Obviously, the choice of examples shown in the prompt may additionally impact the evaluation performance. Nevertheless, the finally proposed evaluation prompt seems to well suit both domains: synthetic and real-world, and should potentially be effective in other similar datasets and solving strategies.
@@
0.93 0.9 0.94 0.9 0.96 0.99
Figure 5: Synthetic BP #85. This is the only BP instance for which the voting-based evaluation of MLLM solutions differed from our manual evaluation. Left: Three parts. Right: Five parts. GPT-4o answer: Left: All images are composed of exactly three lines. Right: All images are composed of more than three lines. Voting marked it as incorrect, whereas in manual evaluation it was marked as correct. The first difference lies in the meaning of the words lines and parts, which, in this visual context, seems identical. The second difference stems from the number of the parts on the right side of the problem. The answer seems to be correct, as obviously, five is more than three. However, one could argue that the answer is incomplete, as each of the squares on the right side clearly depicts exactly five parts.
@@
Appendix C Bongard-RWR dataset
Bongard-RWR dataset developed in this work is attached in the technical appendix and will also be released for the reserach community under the MIT license. The dataset generation algorithm is presented in Algorithm 1 using notation introduced in Section 4.1.
1:A set of synthetic concepts $Y$
2:A set of generated instances ${\cal RWR}$
@@ Appendix C Bongard-RWR dataset
Furthermore, Fig. 8 provides additional examples of the proposed Bongard-RWR dataset. Each subfigure presents a comparison between the synthetic Bongard problem and its respective real-world translation in Bongard-RWR. Examples 8(a) and 8(b) were translated automatically, whereas 8(c) and 8(d) were constructed fully manually, including building an appropriate scene and taking a picture. Additionally, Fig. 7 shows a particular approach taken when translating a given synthetic BP to its Bongard-RWR counterpart (problems not translated and those rejected after translation are combined into one category).
@@
(a) Synthetic BP #8 with its automatically translated Bongard-RWR version. Left: Ends of the curve are far apart. Right: Ends of the curve are close together.
(b) Synthetic BP #10 with its automatically translated Bongard-RWR version. Left: Triangles. Right: Quadrangles.
(c) Synthetic BP #41 with it manually constructed Bongard-RWR version. Left: Outline circles on one straight line. Right: Outline circles not on one straight line.
(d) Synthetic BP #47 with its manually constructed Bongard-RWR version. Left: Triangle on top of the circle. Right: Circle on top of the triangle.
@@
Appendix D Coverage of Bongard-RWR instances
Even though the final results of individual models and strategies solving Bongard-RWR are somewhat unsatisfactory, especially in the case of open language response generation, it is worth to examine the degree of overlap of correct answers across the tested models. On the one hand, the results of ground-truth classification and model’s disagreement on the solution evaluation clearly confirm inability of any single model to solving all problems from the Bongard-RWR datatset. On the other hand, it is likely that the overlap is not complete, and it is posible to expand the solution coverage by appropriate mixture-of-experts approach. This is indeed the case in our experiments, as illustrated in Figures 9 - 16. While single proprietary MLLMs solved between $9$ and $14$ instances, the number of instances solved by any MLLM equaled $23$. We leave exploration of this path for future research.
@@ Appendix D Coverage of Bongard-RWR instances
Figure 16: Final summary of all experiments on the Bongard-RWR dataset. The colormap aggregates all problems solved by any tested model using any generation prompting strategy. Overall, the models collectively managed to solve 23 problems.
Appendix E Comparison of Synthetic Bongard vs. Bongard-RWR Results
Our research shows that all of the tested models have difficulty solving synthetic concepts when applied to real-world images. Comparing the results for both datasets (see Figures 10 and 16) we identified some discrepancies. Four problems that remained unsolved in the synthetic BPs were successfully solved in the real-world domain of the Bongard-RWR dataset: #56, #87, #88, and #98. However, three of these problems differ slightly from their synthetic counterparts. The images in #56 from Bongard-RWR feature a variety of colors instead of the usual black-and-white figures. Furthermore, in real-world version of problem #87 more images feature disjoint elements instead of multi-part objects, which may have nudged the model toward the correct answer. Additionally, in problem #98, the figures are shown against a hatched texture, which was not accounted for in the real-world translation.
Conversely, 22 problems that were solved in the synthetic BPs were not solved in the Bongard-RWR dataset. In most cases, the models focused on general associations that did not apply to all the images. For example, the concept presented in problem #3 is: ”LEFT: Outline figures, RIGHT: Solid figures.” Claude 3.5 Sonnet, using the Descriptive strategy, responded: ”LEFT: focuses on practical, everyday objects or scenes, RIGHT: emphasizes aesthetic, artistic, or decorative elements that serve more for visual appeal than utility”. Nevertheless, both sides feature a red coffee cup with a saucer, which matches both generated descriptions. The key difference lies in the color of the saucer’s rim, which determines whether it should be considered as an outline or a solid figure.
@@
Figure 17: Synthetic BP #56 with its automatically translated Bongard-RWR version. Correct answer: ”Left: All figures of the same color. Right: Figures of different colors”. Provided answers belong to Gemini 1.5 Pro using the Contrastive-iterative strategy, as it was the only combination that solved this Bongard-RWR instance correctly.
@@
Figure 18: Synthetic BP #87 with its manually translated Bongard-RWR version. Correct answer: ”Left: Four parts. Right: Five parts”. Provided answers belong to Claude 3.5 Sonnet using the Descriptive strategy, as it was the only combination that solved this Bongard-RWR instance correctly.
@@
Figure 19: Synthetic BP #88 with its manually translated Bongard-RWR version. Correct answer: ”Left: Three parts. Right: Five parts”. Provided answers belong to GPT-4o using the Direct strategy. Overall this problem was solved by three models and four strategies.
@@
Figure 20: Synthetic BP #98 with its manually corrected Bongard-RWR version. Correct answer: ”Left: Three parts. Right: Five parts”. Provided answers belong to GPT-4 Turbo using the Descriptive-direct strategy. Overall this problem was solved by two models using two different strategies.

### v2

@@ Appendix E Evaluation of MLLMs Answers
Open-ended characteristics of BPs stemming from a textual form of an answer, and the number of considered models ($8$), generation strategies ($7$), datasets ($4$), and BP instances per dataset ($60$ in Bongard-RWR and $100$ in the remaining cases) require the use of an automated NLP-based evaluation of the model’s answers. For this task we employed MLLMs with a specially designed prompt. The initial version of the evaluation prompt (see Prompt E.1) was intentionally relatively simple – a model received an answer to be evaluated as well as the ground-truth labels, and was requested to output a binary yes/no answer. This prompt formulation turned out to be too simplistic. While the level of agreement between all models was relatively high ($87\%$ of responses were rated unanimously by all models), as illustrated in Fig. 11(a), manual inspection of the selected answers revealed that the assessment was generally too optimistic and relatively many evaluations wrongly pointed to correct answers.
@@
E.1 Evaluation Prompt Optimization
Due to the above evaluation disagreement, we made an attempt to optimize the prompt based on the GPT-4o solutions for the additional $20$ BPs (#$\#101-\#120$) that were not used in the main experiments. First, following the few-shot prompting technique, we expanded the prompt to include two examples showing a possible logical difference between correct and incorrect answers. Furthermore, we added a sentence which requested a strict logical compliance with the provided labels. However, this refinement appeared to be too strong, as $2$ (out of $4$) models didn’t evaluate any of the solutions as correct.
To impose some flexibility, we changed the word strictly to logically, but this resulted in an increased rate of false-positive evaluations. Finally, we combined these two prompts, obtaining the outcome closest to the manual (our human) evaluation. The final version of our evaluation prompt is listed in Prompt E.1. Additionally, we attempted to attach the image of the evaluated BP instance to each version of our prompt, but this actually confused the models rather than improving their results, so we ultimately abandoned this option and stuck with the fully text-based prompt.
Although the consistency of results regarded as the number of unanimous assessments stayed at the same level ($87\%$) (see Fig. 11(b)), the number of answers rated as correct significantly decreased, which was in accordance with our random manual verification.
@@ E.1 Evaluation Prompt Optimization
[tbp] Prompt 2: Initial prompt used in MLLM answer evaluation. We focused on its clarity and simplicity. You are a logic module designed to provide accurate answers. In a Bongard Problem the objective is to spot the difference between the contents of images located on the two opposite sides of the problem. You are given correct labels of these sides and must decide whether the answer provided by the user is correct and matches with those labels. Answer with ’OK’ or ’WRONG’. LEFT SIDE LABEL: <left_label> RIGHT SIDE LABEL: <right_label> USER ANSWER: <model_answer>
[tbp] Prompt 3: The final version of the evaluation prompt. It is enriched with the few-shot prompting technique and imposes a logical compliance with provided labels. You are a logic module designed to provide accurate answers. In a Bongard Problem the objective is to spot the difference between the contents of images located on the two opposite sides of the problem. You are given correct labels of these sides and must decide whether the answer provided by the user is correct and matches with those labels. Answer with ’OK’ or ’WRONG’. The user’s answer has to strictly logically match the labels, as shown in examples. FIRST EXAMPLE: LEFT SIDE LABEL: All shapes are small. RIGHT SIDE LABEL: All shapes are big. USER ANSWER: On the left side, one of the shapes is small. On the right side, all of the shapes are big. EVALUATION: WRONG END OF FIRST EXAMPLE. SECOND EXAMPLE: LEFT SIDE LABEL: All shapes are small. RIGHT SIDE LABEL: All shapes are big. USER ANSWER: On the left side, all of the shapes are small. On the right side, all of the shapes are big. EVALUATION: OK END OF SECOND EXAMPLE. LEFT SIDE LABEL: <left_label> RIGHT SIDE LABEL: <right_label> USER ANSWER: <model_answer>
E.2 Manual Verification of the Evaluation Performance of MLLMs
@@
In order to finally assess the efficacy of Prompt E.1 we manually checked the models’ evaluation performance on the $100$ BPs solved by GPT-4o using the Descriptive strategy. As shown in Table 4 in the main paper, all proprietary models achieved better scores on incorrect labels classification. For this reason, we decided to manually verify only the problems evaluated as correct by at least one MLLM, assuming that those incorrect are generally evaluated properly. The comparison between the evaluation performance of the initial and final prompts and our manual evaluation is presented in Table 5. All models denotes evaluations where a solution is marked as correct only if all models evaluate it as correct. Similarly, any model refers to the cases where a solution is marked as correct if at least one model evaluates it as correct. Voting refers to the hard-voting scheme described in section E.1. It is important to observe that the chosen voting evaluation method differed from the manual evaluation in only one specific problem, which is depicted in Fig. 12.
@@
In addition, we checked the performance of our enhanced evaluation prompt on $20$ new, not used in other experiments, manually evaluated Bongard-OpenWorld problems solved by GPT-4o using the Descriptive strategy. Again, the use of Prompt E.1 visibly increased the consensus with manual evaluation (see Table 6). The difference between our manual evaluation and the voting scheme occurred only in $2$ problem solutions whose correctness is disputable (see Fig. 13).
@@
While the choice of examples shown in the prompt may in some cases impact the evaluation performance, the finally proposed evaluation prompt seems to be well suited to both domains: synthetic and real-world, and should potentially be effective in other similar datasets and solving strategies.
@@
$0.93$ $0.90$ $0.94$ $0.90$ $0.96$ $0.99$
Figure 12: Synthetic BP #$\#85$. This is the only BP instance for which the voting-based evaluation of MLLM solutions differed from our manual evaluation. Left: Three parts. Right: Five parts. GPT-4o answer: Left: All images are composed of exactly three lines. Right: All images are composed of more than three lines. Voting marked it as incorrect, whereas in manual evaluation it was marked as correct. The first difference lies in the meaning of the words lines and parts, which, in this visual context, seems identical. The second difference stems from the number of the parts on the right side of the problem. The answer seems to be correct, as obviously, five is more than three. However, one could argue that the answer is incomplete, as each of the squares on the right side clearly depicts exactly five parts.
@@
Appendix F Bongard-RWR Dataset
The dataset generation algorithm is presented in Algorithm 1 using notation introduced in Section 4.
1:  Input: A set of synthetic concepts $Y$
2:  Output: A set of generated instances ${\cal RWR}$
@@ Appendix F Bongard-RWR Dataset
Furthermore, Fig. 15 provides additional examples of the proposed Bongard-RWR dataset. Each subfigure presents a comparison between the synthetic Bongard problem and its respective real-world translation in Bongard-RWR. Examples 15(a) and 15(b) were translated automatically, whereas 15(c) and 15(d) were constructed fully manually, including building an appropriate scene and taking a picture. Additionally, Fig. 14 shows a particular approach taken when translating a given synthetic BP to its Bongard-RWR counterpart (problems not translated and those rejected after translation are combined into one category).
@@
(a) Synthetic BP #$\#8$ with its automatically translated Bongard-RWR version. Left: Figures on the right side. Right: Figures on the left side.
(b) Synthetic BP #$\#10$ with its automatically translated Bongard-RWR version. Left: Triangles. Right: Quadrangles.
(c) Synthetic BP #$\#41$ with it manually constructed Bongard-RWR version. Left: Outline circles on one straight line. Right: Outline circles not on one straight line.
(d) Synthetic BP #$\#47$ with its manually constructed Bongard-RWR version. Left: Triangle on top of the circle. Right: Circle on top of the triangle.
@@
Appendix G Coverage of Bongard-RWR Instances
Even though the final results of individual models and strategies solving Bongard-RWR are somewhat unsatisfactory, especially in the case of open language response generation, it is worth to examine the degree of overlap of correct answers across the tested models. On the one hand, the results of ground-truth classification and model’s disagreement on the solution evaluation clearly confirm inability of any single model to solving all problems from the Bongard-RWR dataset. On the other hand, it is likely that the overlap is not complete, and it is possible to expand the solution coverage by appropriate mixture-of-experts approach. This is indeed the case in our experiments, as illustrated in Figs. 1623. While single proprietary MLLMs solved between $9$ and $14$ instances, the number of instances solved by any MLLM equaled $23$. We leave exploration of this path for future research.
@@ Appendix G Coverage of Bongard-RWR Instances
Figure 23: Final summary of all experiments on the Bongard-RWR dataset. The colormap aggregates all problems solved by any tested model using any generation prompting strategy. Overall, the models collectively managed to solve $22$ problems.
Appendix H Comparison of Synthetic Bongard vs. Bongard-RWR results
Our research shows that all of the tested models have difficulty solving synthetic concepts when applied to real-world images. Comparing the results for both datasets (see Figs. 17 and 23) we identified some discrepancies. Four problems that remained unsolved in the synthetic BPs were successfully solved in the real-world domain of the Bongard-RWR dataset: #$\#56$, #$\#87$, #$\#88$, and #$\#98$. However, three of these problems differ slightly from their synthetic counterparts. The images in #$\#56$ from Bongard-RWR feature a variety of colors instead of the usual black-and-white figures. Furthermore, in real-world version of problem #$\#87$ more images feature disjoint elements instead of multi-part objects, which may have nudged the model toward the correct answer. Additionally, in problem #$\#98$, the figures are shown against a hatched texture, which was not accounted for in the real-world translation.
Conversely, $22$ problems that were solved in the synthetic BPs were not solved in the Bongard-RWR dataset. In most cases, the models focused on general associations that did not apply to all the images. For example, the concept presented in problem #$\#3$ is: ”LEFT: Outline figures, RIGHT: Solid figures.” Claude 3.5 Sonnet, using the Descriptive strategy, responded: ”LEFT: focuses on practical, everyday objects or scenes, RIGHT: emphasizes aesthetic, artistic, or decorative elements that serve more for visual appeal than utility”. Nevertheless, both sides feature a red coffee cup with a saucer, which matches both generated descriptions. The key difference lies in the color of the saucer’s rim, which determines whether it should be considered as an outline or a solid figure.
@@
Figure 25: Synthetic BP #$\#87$ with its manually translated Bongard-RWR version. Correct answer: ”Left: Four parts. Right: Five parts”. Provided answers belong to Claude 3.5 Sonnet using the Descriptive strategy, as it was the only combination that solved this Bongard-RWR instance correctly.
@@
Figure 26: Synthetic BP #$\#88$ with its manually translated Bongard-RWR version. Correct answer: ”Left: Three parts. Right: Five parts”. Provided answers belong to GPT-4o using the Direct strategy. Overall this problem was solved by three models and four strategies.
@@
Figure 27: Synthetic BP #$\#98$ with its manually corrected Bongard-RWR version. Correct answer: ”Left: Three parts. Right: Five parts”. Provided answers belong to GPT-4 Turbo using the Descriptive-direct strategy. Overall this problem was solved by two models using two different strategies.
Appendix I MLLM Prompts
I.1 Prompts for Bongard-RWR Generation
At first, Prompt I.1 was employed to generate real-world representations of synthetic concepts. After downloading the images corresponding to these textual descriptions, Prompt I.1 was used to select those images that correctly represent given concept translation. In addition to the left and right concepts, we also provided prompts briefly explaining the context that the image should match. These prompts were generated during the translation stage of our algorithm (see the fourth line in Algorithm 1).
[tbp] Prompt 4: Initial concept-describing prompt used in construction of Bongard-RWR. Your goal is to translate a comparison concept from the geometric domain to the real-world domain. Your translations should be expressible as images. Example: Geometric domain: triangles vs squares {   "left": {     "concept": "pyramids"   },   "right": {     "concept": "rectangular buildings"   } } Give <number> unique translations for the following concept as a raw JSON array of objects (same as in the example above). <concept>
[tbp] Prompt 5: Prompt used for the selection of proper images for a translated concept. You translated a concept comparison from geometric domain to the real-world domain as follows: Geometric domain: <left_geometric_concept> vs <right_geometric_concept> Real world domain: {     "left": {         "concept": <left_concept>,         "prompt": <left_concept_description>     },     "right": {         "concept": <right_concept>,         "prompt": <right_concept_description>     } } Now, you need to check if the queried image matches your translation and provides enough information to distinguish it from the other concept. Don’t focus too much on the prompt. It’s just a hint for you to understand the concept better. Provided image represents <side_concept> Give your answer in the following format: EVALUATION: OK EXPLANATION: <here you can provide additional information> or EVALUATION: REJECTED EXPLANATION: <here you can provide additional information>
I.2 Prompt Describing the Bongard Problem
Prompt I.2 describing the BP task has been placed at the beginning of each solving strategy introduced in Section 3.1.
[tbp] Prompt 6: Prompt explaining Bongard problems to an MLLM. A Bongard Problem is composed of left and right sides separated by a line. Each side contains six images. All images belonging to one side present a common concept, which is lacking in all images from the other side. The goal is to describe the rule that fits all images on the left side, but none on the right, and, conversely, the rule that fits all images on the right side, but none on the left. The description of the rule should be simple and concise. Example 1: All shapes on left are small. All shapes on right are big. Example 2: The left side contains circles. The right side contains triangles.
I.3 Prompts for Classification Strategies
Prompt I.3 was used to assess solution correctness (see Section 3.2). Prompt I.3 was used to assign images to sides (see Section 3.3).
[tbp] Prompt 7: Prompt used for images to side classification (see Fig. 3c). Two examples were provided to not bias the results of the model. You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to classify two test images to the corresponding side of the Bongard Problem, LEFT or RIGHT. Each image belongs to exactly one class. The test images belong to different classes. The images are always provided correctly. Respond only to the specific request. Respond in json using the following format. FIRST EXAMPLE: Left images: <small shapes> Right images: <big shapes> First test image: <small shape> Second test image: <big shape> Response: {     "first": {         "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.",         "concept": "small vs big",         "answer": "LEFT"     },     "second": {         "explanation": "The test image shows a big shape, similarly as all images on right. The images on left, on the other hand, feature small shapes.",         "concept": "small vs big",         "answer": "RIGHT"     } } END OF FIRST EXAMPLE SECOND EXAMPLE: Left images: <circles> Right images: <triangles> First test image: <triangle> Second test image: <circle> Response: {     "first": {         "explanation": "The test image shows a triangle, which matches all images on right. In contrast, the left side images feature circles.",         "concept": "circles vs triangles",         "answer": "RIGHT"     },     "second": {         "explanation": "The test image shows a circle, which matches all images on left. Conversely, the right side images feature triangles.",         "concept": "circles vs triangles",         "answer": "LEFT"     } } END OF SECOND EXAMPLE
[tbp] Prompt 8: Prompt used for the solution correctness assessment (see Fig. 3b). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to evaluate the correctness of the provided answer to the given Bongard Problem. All images are provided correctly. Do not explain the answer, just evaluate it. Respond ’OK’ if the answer is correct, otherwise respond ’WRONG’. User answer: <user_answer>
I.4 Prompts for Natural Language Answer Generation Strategies
Prompts I.4I.4 were used for natural language answer generation (see Section 3.1).
[tbp] Prompt 9: Prompt used for the Direct strategy. (see Fig. 2a). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem. What is the difference between the two sides of the problem?
[tbp] Prompt 10: Prompt used to obtain the image descriptions in the Descriptive strategy (see Fig. 2b). The provided image is a part of an abstract visual reasoning problem. Describe all crucial properties of the image. Your description should be as concise as possible. Focus on the most important details. The image is provided correctly. Respond only with descriptions.
[tbp] Prompt 11: Prompt used for the Descriptive and Descriptive-direct strategies (see Fig. 2b). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem using descriptions of its images. LEFT IMAGES: <left_descriptions> RIGHT IMAGES: <right_descriptions> What is the difference between the two sides of the problem?
[tbp] Prompt 12: Prompt used to obtain the image descriptions in the Descriptive-iterative strategy (see Fig. 2c). After the last image, we used the prompt: “That was the last image. Now provide your final answer.” You’ll receive a sequence of images that are a part of a single side of a Bongard Problem. The images will be provided one by one. Your goal is to find a common concept presented in all images. Your description should be as concise as possible. Focus on the most important details. Try to enhance the description of the concept after each image. The image is always provided correctly. Respond only to the specific request. The first image will be provided in the next message.
[tbp] Prompt 13: Prompt used for the Descriptive-iterative strategy (see Fig. 2c). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem using descriptions of two sides of the problem. LEFT SIDE DESCRIPTION: <left_description> RIGHT SIDE DESCRIPTION: <right_description> What is the difference between the two sides of the problem?
[tbp] Prompt 14: Prompt used to obtain the comparison between the left and right image in the Contrastive strategy (see Fig. 2d). After the last image, we used the prompt: “That was the last image. Now provide your final answer.” You are given two images extracted from the left and right side of a Bongard Problem, respectively. Your goal is to compare the images. Your comparison should be as concise as possible.
[tbp] Prompt 15: Prompt used for the Contrastive and Contrastive-direct strategies (see Fig. 2d). You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem using comparisons between pairs of images. Each pair contains one image from the left and one from the right side of the problem. COMPARISONS: <comparisons> What is the difference between the two sides of the problem?
[tbp] Prompt 16: Prompt used for the Contrastive-iterative strategy (see Fig. 2e). After the last pair of images, we used the prompt: “It was the last pair of images. What is the difference between the two sides of the problem?” You are a vision understanding module designed to provide short, clear and accurate answers. Your goal is to solve the provided Bongard Problem. You’ll receive a sequence of image pairs. Each pair contains one image from the left and one from the right side of the problem. In each step compare the two images and refine the definitions of concepts that describe left and right sides of the problem. Your description should be as concise as possible. Focus on the most important details. The first pair will be provided in the next message.